suppose u is a subspace of r^4 defined by u=span((1,2,3,-4),(-5,4,3,2)). find an orthonormal basis of u and an orthonormal basis of u^Ï€


1
Expert's answer
2022-01-31T15:54:03-0500

Let "v_1=(1,2,3,-4)", "v_2=(-5,4,3,2)", "v_3=(0,0,1,0)", "v_4=(0,0,0,1)". This is a basis of the vector space "{\mathbb R}^4", "v_1" and "v_2" form a basis of "U". Let's apply the Gram-Schmidt's method of orthogonalization to this vector system.


Let "w_1=v_1=(1,2,3,-4)", then:

"\langle{w_1,w_1}\rangle=1^2+2^2+3^2+(-4)^2=30".

"\langle{v_2,w_1}\rangle=-5\cdot 1+4\cdot 2+3\cdot 3+2\cdot(-4)=4".

"\langle{v_3,w_1}\rangle=3".

"\langle{v_4,w_1}\rangle=-4".


Calculate "v_2-\frac{\langle{v_2,w_1}\rangle}{\langle{w_1,w_1}\rangle} w_1=v_2-\frac{4}{30}w_1="

"=(-5,4,3,2)-\frac{2}{15}(1,2,3,-4)=\frac{1}{15}(-77,56,39,38)".


Let "w_2=(-77,56,39,38)", then:

"\langle{w_1,w_2}\rangle=0".

"\langle{w_2,w_2}\rangle=(-77)^2+56^2+39^2+38^2=12030"

"\langle{v_3,w_2}\rangle=39".

"\langle{v_4,w_2}\rangle=38".


Calculate "v_3-\frac{\langle{v_3,w_1}\rangle}{\langle{w_1,w_1}\rangle} w_1-\frac{\langle{v_3,w_2}\rangle}{\langle{w_2,w_2}\rangle} w_2=v_3-\frac{3}{30}w_1-\frac{39}{12030}w_2"

"=\frac{1}{4010}(4010v_3-401w_1-13w_2)=\frac{1}{4010}(600,-1530,2300,1110)"

"=\frac{1}{401}(60,-153,230,111)".


Let "w_3=(60,-153,230,111)". Then

"\langle{w_1,w_3}\rangle=0"

"\langle{w_2,w_3}\rangle=0"

"\langle{w_3,w_3}\rangle=60^2+(-153)^2+230^2+111^2=92230"

"\langle{v_4,w_3}\rangle=111".


Calculate "v_4-\frac{\langle{v_4,w_1}\rangle}{\langle{w_1,w_1}\rangle} w_1-\frac{\langle{v_4,w_2}\rangle}{\langle{w_2,w_2}\rangle} w_2-\frac{\langle{v_4,w_3}\rangle}{\langle{w_3,w_3}\rangle} w_3"

"=v_4+\frac{4}{30}w_1-\frac{38}{12030}w_2-\frac{111}{92230}w_3"

"=\frac{1}{276690}(276690v_4+27669w_1-874w_2-333w_3)"

"=\frac{1}{276690}(74987,57343,-27669,95839
)".


Let "w_4=(74987,57343,-27669,95839)". Then

"\langle{w_1,w_4}\rangle=0"

"\langle{w_2,w_4}\rangle=0"

"\langle{w_3,w_4}\rangle=0"

"\langle{w_4,w_4}\rangle=18861957300".


We have constructed an orthogonal basis "w_1,\,w_2,\,w_3,\,w_4" of "{\mathbb R}^4". The vectors "w_1,\,w_2" form a basis of U, because "w_1,w_2\in Span\{v_1,v_2\}=U" and "\dim U=2".

Since the vectors "w_3,\,w_4" are orthogonal to "w_1,w_2", they form a basis of "U^{\perp}".

To obtain orthonormal bases, one should normalize these vectors and take "\frac{1}{\sqrt{30}}w_1", "\frac{1}{\sqrt{12030}}w_2", "\frac{1}{\sqrt{92230}}w_3" and "\frac{1}{\sqrt{18861957300}}w_4" instead of them.